### v2

# History of $N(1680)$

Abstract
This paper describes my personal appreciation for some of the great research achievements of Mitya Diakonov, Vitya Petrov, and Maxim Polyakov and how my own research career has followed the paths they opened. Among the topics where they have been the most influential have been the pursuit and study of the exotic pentaquark. The search for exotics may require a complementary approach, such as experimental and theoretical activity. Here, I would like to focus on a story of $N(1680)$ in which I was involved.
Dedication
This paper is dedicated to the memory of Mitya, Vitya, and Maxim. Perhaps now that they are no longer with us, we can better appreciate their scientific achievements.
1 Prehistory
I don’t remember Maxim well from the Leningrad Nuclear Physics Institute (LNPI and PNPI now) Winter School on Nuclear and Particle Physics times in the 1970s when he was a student. But somewhere in 1997 or so (the time of the famous $\overline{\textbf{10}}$ paper [1] came up), Dick Arndt forwarded me a letter from Maxim asking what the SAID kaon-nucleon Partial-Wave Analysis (PWA) [2] gives for the exotic $S=+1$ baryon. Dick was very surprised, since the analysis gave nothing which may fit such a narrow exotic state (all four resonances in standard PWA are too heavy $\mathrm{ReW}>1790~{}\mathrm{MeV}$ and too broad $\mathrm{|ImW|}>95~{}\mathrm{MeV}$, then so do not fit the predicted $\Theta^{+}$ in Ref. [1]). From that moment on, my friendship and long-term cooperation with Maxim began. To solve a PWA problem, we came up with a modified PWA to look for the narrow resonance [3].
I met Mitya for the first time at one of the LNPI Winter Schools in the mid 1970s. It is known that the mass difference in multiplets is determined by the phenomenological Gell-Mann-Okubo mass formula, which depends on the nucleon sigma term. Our SAID sigma term [4] was selected for this purpose [1]. In 2005 or so, Mitya gave a talk at a JLab seminar and Dick Arndt, who was sitting next to me, asked where this boy got such English. My answer was simple - his mother (Nina Yakovlevna) translated a poem by John Keats from English into Russian. Dick was satisfied with my answer.
I also met Vitya for the first time at one of the LNPI Winter Schools in the 1970s, but there is a background. In 1968, I measured the efficiency of neutron detection with large scintillation counters, using a beam of neutrons from the Neutron Generator (that was my MS project). At the same time, a physicist and an engineer studied the reaction of transistors to neutrons using the same source of fast neutrons. The physicist’s name was Tamara Markovna. One day, a strange man appeared looking for his wife. Somebody told me who he was, Yuri Viktorovich Petrov, a well-known theoretician. That’s how I found out Tamara Markovna’s last name (they were Vitya’s parents), and Vitya was maybe 13 years old at the time.
Obviously, the Instanton Vacuum model is a very powerful tool [5]. Before the $\overline{\textbf{10}}$ tasks came up, our friends focused on the nucleon-nucleon interaction and, in particular, on the dibaryon (see, for instance, Ref. [6]). In 1997, my postdoc, Serezha Pavlenko, started to work with Vitya and Mitya. He was a brilliant young experimentalist and phenomenologist. In a short time, Serezha did many calculations and Mitya and Vitya had a high respect for Serezha. Unfortunately, a stroke killed him. In 1998, he was 24 years old.
2 Introduction
QCD gives rise to the hadron spectrum [7] and many $q\bar{q}$ and $qqq$ have been observed [8]. However, $q\bar{q}q\bar{q}$ and $qqqq\bar{q}$ and other many-quark states are not forbidden but have not yet been observed. Recently, the LHCb Collaboration claimed evidence for four hidden-charm $qqqc\bar{c}$ states near open-charm decay thresholds for $\Sigma_{c}^{+}\bar{D}^{0}$ and $\Sigma_{c}^{+}\bar{D}^{\ast 0}$ in $\Lambda_{b}\to J/\psi pK^{-}$ decay [9]. Nevertheless, although there is no doubt about the LHCb observations, it is not clear whether they are compact multi-quark or hadonic molecular states [10].
In the light quark sector, Diakonov, Petrov, and Polyakov proposed a clearly exotic and narrow $\Theta^{+}(uudd\bar{s})$ state [11], yet to be unequivocally observed and identified. The original name of the pentaquark lying at the apex of $\overline{\textbf{10}}$ was $Z^{+}$, then following Mitya’s suggestion, we called it $\Theta^{+}$ now (see Appendix). Using the Chiral Quark Soliton model [11], this particle along with other members of $\overline{\textbf{{10}}}$ has been proposed with a mass of $M_{\Theta^{+}}=1.53~{}\mathrm{GeV}$ and a width less than $15~{}\mathrm{MeV}$ [1] (Fig. 1). Due to relatively low mass and simple decay channels to $K^{+}n$ or $K^{0}p$, it has attracted attention of many experiments at different facilities worldwide (the Citation Index (CI) on this renowned paper [1] is about 1000).
Figure 1: The suggested anti-decuplet of baryons [1]. The three corners of the triangle (in red) are exotic, which means that their quantum numbers require more than three quarks.
The first experimental evidence for $\Theta^{+}$ came from the LEPS Collaboration at SPring-8 (investigating reaction $\gamma n\to K^{+}K^{-}n$ on 12C) [12] and the DIANA Collaboration from ITEP (investigating reaction $K^{+}n\to K^{0}p$ using data from the Xe bubble chamber) [13].
The standard PWA (by construction) tends to miss a narrow resonance with $\Gamma_{R}<20~{}\mathrm{MeV}$ or so [3]. Modified PWA assumes the existence of a narrow resonance by inserting a Breit-Wigner (BW) structure to one of the partial amplitude and refitting the whole database. Such an analysis of the effect of the scanning mass $M_{R}$, full $\Gamma_{R}$ and partial width of the narrow resonance for each partial wave allows us to get a picture. The worse description of the full database that includes resonance with the corresponding $M_{R}$ and $\Gamma_{R}$ (in terms of $\Delta\chi^{2}$) does not support our hypothesis. The better description may result that the resonance may exist, or the effect can be due to various corrections (e.g. thresholds), or both possibilities can contribute. Some additional checks are necessary: true resonance should provide the effect only in a particular partial wave, while a non-resonance source may show similar effects in various partial waves.
Using this kind of modified PWA, Arndt and co-workers performed a reanalysis of the existing $KN$ database [2] in order to refits $K^{+}N$ observables and consider the effect of a narrow state [3].
It was found that the existence of a $\Theta^{+}$ in the $P_{01}$ ($J^{P}=1/2^{+}$) state, with a mass of $\sim 1545~{}\mathrm{MeV}$ and a width of $\Gamma(\Theta^{+})\leq 0.5~{}\mathrm{MeV}$ or so is possible (Fig. 2).
Figure 2: Change of overall $\chi^{2}$ due to insertion of a resonance into $P_{01}$ for $M_{R}=1520–1560~{}\mathrm{MeV}$ with $\Gamma_{R}=0.3$ and $0.7~{}\mathrm{MeV}$, using KN plane wave approximation [2]. Energies, where measurements exist, are labeled by red vertical bars (for references, see Ref. [2]). The $1535~{}\mathrm{MeV}$ values for $\Delta\chi^{2}$ are off-scale [3]. The lines are given to guide the eye.
3 My Story of $N(1680)$
The mass difference in baryon multiplet is determined by the phenomenological Gell-Mann-Okubo mass formula (mixing is able to shift some masses). Originally, Mitya, Vitya and Maxim claimed $N(1710)$ as a non-strange member of $\overline{\textbf{10}}$, which was a suitable candidate for $N^{\ast}$ [14]. The state $N(1710)$, though listed in the PDG2002 baryon summary table [15] as a three-star resonance, was not seen in the PWA of the $\pi N$ elastic scattering data by the SAID group [16].
Studies that have claimed to see this $N(1710)$ state have given widely varying estimates of its mass (from $1680~{}\mathrm{MeV}$ to $1740~{}\mathrm{MeV}$) and width (from $90~{}\mathrm{MeV}$ to $500~{}\mathrm{MeV}$). Branching ratios have also been given with large uncertainties (10–20% for $N\pi$, 40–90% for $N\pi\pi$, etc.), apart from one that has been presented with greater precision ($6\pm 1\%$ for $N\eta$[15].
Figure 3: Argand plots for the $P$ partial-wave amplitudes for the $\pi N$ case from threshold ($1080~{}\mathrm{MeV}$) to $W=2500~{}\mathrm{MeV}$ [17]. The crosses indicate $50~{}\mathrm{MeV}$ steps in $W$. Solid circles correspond to the BW $W_{R}$.
The $\pi$N PWA by the GW SAID group has shown that above the Roper resonance, $W>1500~{}\mathrm{MeV}$, $\sigma^{tot}\cong 2\sigma^{el}\cong 2\sigma^{inel}$ ($\sigma^{tot}=\sigma^{el}+\sigma^{inel}$). It means that is nearly pure diffraction, inelasticity $\eta\to\infty$ for $P_{11}$ and $S\cong 0$ with $A\cong i/2$. The result is that phase $\delta$ is poorly defined and the $P_{11}$ amplitude is spinning around $ReA=0$ [17] (Fig. 3).
Of course, the non-observation of a broad $N(1710)$ state in $\pi N$ elastic analyses (no pole position, no BW, no speed plot) could be due to a very small $\pi N$ branching ratio. Standard procedure used in PWA may also miss narrow resonances with $\Gamma<20~{}\mathrm{MeV}$ by construction [16]. Therefore, the true unitary partner of $\Theta^{+}$ (if it is different from $N(1710)$ and sufficiently narrow) could have eluded detection.
Arndt and co-workers were considering the $\pi N$ partial wave $P_{11}$, as this is amplitude, which is associated with resonances having $J^{P}=1/2^{+}$ [18]. The character of $\chi^{2}$ changes, $\Delta\chi^{2}$, after inserting a narrow resonance into the partial amplitude with a range of masses, widths, and branching fractions is illustrated in Fig. 4. Negative values of $\Delta\chi^{2}$ emerge more readily near $M_{R}=1680~{}\mathrm{MeV}$ and $1730~{}\mathrm{MeV}$. We see that $\Delta\chi^{2}$ becomes negative only for $\Gamma_{el}=(\Gamma_{el}/\Gamma_{tot})\cdot\Gamma_{tot}$ within the bounds $\Gamma_{el}\leq 0.5~{}(0.3)~{}\mathrm{MeV}$ for $M_{R}=1680~{}(1730)~{}\mathrm{MeV}$ (Fig. 4). The available data cannot reliably discriminate values of $\Gamma_{el}$ below these bounds.
[…]
The expected decay properties of $N(1680)$ are essentially model-dependent and the Chiral Quark Soliton approach with violated SU(3)F [mixing $N_{\overline{\textbf{{10}}}}-N_{\textbf{8}}$] gives for $M_{N^{\ast}}=1680~{}(1730)~{}\mathrm{MeV}$ [18]:
  • $\Gamma(N^{\ast}\to\pi\Delta)\approx 2.8~{}(3.5)~{}\mathrm{MeV}$.
  • $\Gamma(N^{\ast}\to K\Lambda)\approx 0.70~{}(1.56)~{}\mathrm{MeV}$.
  • $\Gamma(N^{\ast}\to\eta N)\approx 2~{}\mathrm{MeV}$.
  • $\Gamma(N^{\ast}\to\pi N)\approx 2.1~{}(2.3)~{}\mathrm{MeV}$.
  • $\Gamma(N^{\ast}\to\mathrm{tot})\approx 10~{}\mathrm{MeV}$.
Then, the GRAAL Collaboration claimed evidence for a narrow ($10~{}\mathrm{MeV}$) resonance state with a mass of $1675~{}\mathrm{MeV}$ in the reaction $\gamma n\to\eta n$
(Fig. 5 (left)) [19]. BTW, in the case of a “neutron” target, Final State Interaction (FSI) is a problem. The effect is absent in the case of $\gamma p\to\eta p$ (Fig. 5 (right)) and agrees with the prediction of the Chiral Quark Solution approach [20]. In exact SU(3)F, the transition magnetic moment $\mu(p^{\ast}\to p)$ should vanish, since the U-spins are 3/2 for $p^{\ast}$, 1/2 for $p$, and 0 for the photon. With violation of SU(3)F, this transition moment does not vanish but is still much smaller than $\mu(n^{\ast}\to n)$ [21].
Figure 5: Quasi-free $\eta n$ (left) and free $\eta p$ (right) photoproduction cross sections (dark circles) [19]. The dashed area (left plot) shows the simulated contribution of a narrow state at $W=1.675~{}\mathrm{MeV}$. The solid line (right plot) indicates the solution of E429 of the SAID $\gamma p\to\eta p$ PWA [22].
The GW SAID group performed a coupled channel analysis of the $\pi N$ system including $\pi^{\pm}p$ elastic scattering and $\pi^{-}p\to\pi^{0}n$ with $\pi^{-}p\to\eta n$ [16]. The attractive factor for the reaction $\pi^{-}p\to\eta n$ is that it plays a role in isospin filtering. Unfortunately, the $\pi^{-}p\to\eta n$ data above $800~{}\mathrm{MeV}$ ($W=1630~{}\mathrm{MeV}$) are not reliable for PWA (see, for instance, Fig. 6 (left)). Most of NIMROD data do not satisfy requirements (systematics is 10% or more, momentum uncertainties are up to $100~{}\mathrm{MeV/c}$, and so on). That is the main reasons why the SAID coupled channel analysis had no chance to look for $N(1680)$ using reaction data $\pi^{-}p\to\eta n$ (Fig. 6 (right)).
Figure 6: Left: Excitation function for differential cross sections for the reaction $\pi^{-}p\to\eta n$. Data came from BNL, RHEL, and Saclay. Solid blue line from SAID PWA (WI08 solution) [24]. Right: Total cross sections for the reaction $\pi^{-}p\to\eta n$. Red solid line from SAID PWA [24]. All data are available at Ref. [23].
EPECUR Collaboration at ITEP performed high quality measurements of the differential cross sections for the elastic $\pi^{\pm}p$ over an energy range of $p=800-1300~{}\mathrm{MeV/c}$ ($W=1.55-1.83~{}\mathrm{GeV}$) and for center-of-mass angles from 40 to $120~{}\mathrm{deg}$ [25]. In total, approximately 10,000 new data points have been obtained. These data have been produced with a momentum resolution of $\sim 1~{}\mathrm{MeV/c}$ and with $\sim 1$% statistical uncertainties.
Figure 7: Comparison of the measured $\pi$N data [23] and calculated $\pi^{-}p\to\eta n$ total cross sections. The red solid curve presents the calculations from the work [26]. The black dashed curve indicates the contribution of the $S$ wave.
Following the analysis [18], $\Gamma(\pi N)\sim 0.5~{}\mathrm{MeV}$, which stimulated us to look for $N(1680)$ using high-quality EPECUR elastic $\pi^{\pm}p$ scattering data [26]. The bottom line is that two narrow structures observed in elastic $\pi^{-}p$ scattering can be explained by a combination of threshold effects and two narrow resonances, $S_{11}(1686)$ and $P_{11}(1720)$. Fig. 7 presents the predictions for the $\pi^{-}p\to\eta n$ total cross sections. This model-dependent analysis shows that $S_{11}(1686)$ cannot be a member of $\overline{\textbf{10}}$, while $P_{11}(1720)$ may have a link to the results of Arndt et al. [18].
The new experimental proposal P102 for J-PARC addresses a measurement of the cross section of reaction $\pi^{-}p\to\eta n$ at $p=850-1200~{}\mathrm{MeV/c}$ with a CsI (Tl) calorimeter [27]. The new data will help us solve the puzzle of $N(1680)$.
4 Conclusion
What would Mitya, Vitya, and Maxim make of the current situation in particle physics? On the one hand, they would be justifiably proud of the robust successes of an instanton vacuum approach, which did so much to place it on a firm footing. On the other hand, some of the problems that concerned them remain unresolved, notably the exotic $\Theta^{+}$ and its partners.
Actually, twenty years ago, in 2003, two experimental groups, LEPS and DIANA, announced the observation of a light and narrow exotic baryon with mass in the vicinity of the $1540~{}\mathrm{MeV}$, which was predicted by our friends. The history of this discovery and its theoretical interpretations can be found in the recent review [28].
Acknowledgment
This work was supported in part by the U. S. Department of Energy, Office of Science, Office of Nuclear Physics, under Award No. DE–SC0016583.
Appendix
My archive kept a letter for the name $\Theta^{+}$’ that Mitya sent to the community [29]. The result is that the community accepted Mitya’s name.
From: Diakonov@nordita.dk
To: Takashi Nakano, nakano@rcnp.osaka-u.ac.jp,
Anatoly Dolgolenko,
Elton Smith,
Valeri Koubarovski,
Ken Hicks,
Eugene Pasyuk,
Stepan Stepanyan,
Mark Strikman,
Maxim Polyakov,
Victor Petrov,
Igor Strakovsky,
James Bjorken,
Stan Brodsky,
Lev B. Okun,
Leonid G. Landsberg,
Hiroshi Toki.
Date: Apr 12, 2003, 12:33PM
Subject: exotic baryon name
Dear friends,
Now that, thanks to your efforts, the exotic S=+1 baryon seems to become gradually a reality and not just a fancy dream of a theorist, we have to think how to name the child to be born, hopefully.
The tentative name $Z^{+}$ we all have used is, to my mind, unfortunate. First, I have noticed that the first reaction by people hearing about it is that the new particle has some association with the electroweak Z boson. This association is completely misleading and may create lots of confusion in future and cause us much trouble. Second, the tradition is that baryons are named by capital Greek letters: Delta, Lambda, Sigma, Xi, Omega. (Except the nucleon but one cannot change it.) Third, “Z” is pronounced very differently in different languages; even in English it is “zee” in American English and “zet” in British which is mostly used in Europe and elsewhere. Fourth, and probably most important, the name “Z” was imposed on us by the unsuccessful searches of exotic baryons in the far higher mass region, as summarized in the 1986 edition of Particle Data. This was “prehistoric” time, before the discovery of the Z boson, and in any case nothing has been found worth mentioning in Particle Data Listings since the 1986 edition. The $1540~{}\mathrm{MeV}$ signal has nothing to do with the broad and always suspicious one-star candidates in the $1700~{}\mathrm{MeV}$ region. I regret that we uncritically and unthoughtfully borrowed this queer name and used it in our paper with Petrov and Polyakov. But at that time it was so far from experiment!
I suggest that the name for this baryon should satisfy the following criteria:
1) It must be a capital Greek letter, according to the tradition of naming baryons.
2) It must be distinct from anything used before and carry no associations with bosons.
3) Last but not least, the character must exist in LaTeX.
If you look into the list of upper case Greek letters used in LaTeX you’ll find that there is one and only one character satisfying all criteria, and it is
THETA.
When I discovered it I realized that I kind of liked this name. It is a symmetric and “round” character like the Omega also sitting at the vertex of the big decuplet triangle and, like Omega, alluding to that it is a singlet. If you disregard the historic “tau-theta problem” of the last century’s 50s, this character has been never used in particle physics, as far as I know. It carries no associations and hints that it is something really new, which it certainly is. Last, “Theta” is pronounced more or less in the same way in all languages I know.
Therefore, I suggest that we, experimentalists and theorists, will henceforth call the new candidate for the exotic baryon with strangeness +1 and mass around $1540~{}\mathrm{MeV}$.
$\Theta^{+}$ baryon.
Any objections?
With my best regards,
Dmitri Diakonov.